OpenStudy (anonymous):

let g and h be two functions defined by: see pic below for any real number x. Find g/h(2) and then find all values that are NOT in the domain of g/h. If there is more than one value, separate them with commas.

OpenStudy (anonymous):

\[\frac{ \left( x-1 \right)\left( x-6 \right) }{ -9x+1 }(2)\]

OpenStudy (anonymous):

how do I find values that are not in the domain of g/h(2) @gorv

OpenStudy (anonymous):

@ganeshie8 can you help me

OpenStudy (anonymous):

@***[ISURU]***